Session: Jun 2010 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (G.Maccaferri at Sat May 29 20:41:42 2010) Success:

Onsala (Michael Lindqvist at Sat May 29 20:45:10 2010) Success:
Operator started the wrong experiment so there is some crap data on the first module (from ev018a). We did not loose any data for gb071b though. See log for details.

Yebes (Paco Colomer at Sun May 30 17:18:48 2010) Failure: ANT
The antenna cannot be moved. This experiment has not been observed.

Torun (K. Borkowski at Mon May 31 11:25:46 2010) Success:
No problems but the operator noted two cases of BBC 6 unlocked. The two packs will be send to Socorro shortly, yet today (i.e. Monday, the 31st).

Westerbork (Geert Kuper at Mon May 31 21:00:16 2010) Success:
12 telescopes in the array. No known problems

Effelsberg (Uwe Bach at Wed Jun 02 09:00:30 2010) Success: